AB INITIO CALCULATION OF THE ELECTRONIC TRANSITIONS OF SAMARIUM MONOXIDE

\underline{TIANXIAO YANG} AND RUSSELL M. PITZER, Department of Chemistry, The Ohio State University, 100 W.18th Avenue, Columbus, OH, 43210.

Using relativistic effective core potentials, spin-orbit configuration interaction (SOCI) calculations have been performed to compute the electronic transitions of SmO and SmO^{+}. The ordering and positions of the low-lying states of SmO are shown to be in very good agreement with the experimental measurements^{a}^{,}^{b}and those calculated from ligand field theory.^{c} The SOCI calculations confirmed that the lowest superconfiguration of SmO is Sm^{2+}(4f^{5}6s)O^{2−}. The ground electronic state of SmO is X 0^{−}. The upper state configuration is Sm^{2+}(4f^{5}5pπ)O^{2−}. The electronic transitions of SmO^{+} will also be discussed.

aC. Linton, B. J. Guo, R. S. Rana and J. A. Gray J. Mol. Spectrosc.\underline{126}(370),1987.

bB. J. Guo and C. Linton J. Mol. Spectrosc.\underline{147}(120),1991.

cP. Carette and A. Hocquet J. Mol. Spectrosc.\underline{131}(301),1988.